OpenStudy (anonymous):

The velocity of a diver just before hitting the water is -7.1 m/s, where the minus sign indicates that her motion is directly downward. What is her displacement during the last 0.88 s of the dive?

OpenStudy (anonymous):

x(displacement)=vot+(1/2)at^2 x=-7.1(.88)+(1/20)(9.8*(.88^2)) X=-6.248+3.7946 X=-2.45 X=-2.5

OpenStudy (anonymous):

You have the correct equation and only need to tweek a few things. Check it out...\[\Delta x = v _{i}+\frac{ 1 }{ 2 }a \Delta t ^{2}\] \[\Delta x = -7.1*0.88 + \frac{ 1 }{ 2 }*-9.81*0.88^{2}\]\[\Delta x = -6.248 +(-3.798) = -10.04 m\] Considering significant figures, I would say that the correct answer would be -10.04 m. You neglected to include the "-" sign with the acceleration of gravity and perhaps there was a typo with your 1/2. Right idea though.

OpenStudy (anonymous):

thanks soo much!!